solve each system using substitution
y-2x=3
3x-2y=5

Answer by TutorDelphia(193)

Just by adding 2x to both sides of that first equation we get
y=3+2x
now we can plug in 3+2x in for y
3x-2(3+2x)=5
3x-6-4x=5
-x-6=5
-x=11
x=-11
plug in -11 for x
y-2(-11)=3
y+22=3
y=-19